OpenStudy (anonymous):

Which of the following functions is not a sinusoid? A. y = cos x B. y = sin x C. y = x2 D. None of the above are sinusoids.

OpenStudy (anonymous):

It's C right? I just wasn't sure..

OpenStudy (fifciol):

yes it's C

OpenStudy (anonymous):

Thank you!

OpenStudy (fifciol):

welcome

hartnn (hartnn):

in general, \(A sin (\omega x+\phi)\) is sinusoidal , sin x and cos x fall in this category :)

OpenStudy (anonymous):

Oh! Wait, okay... I see now

hartnn (hartnn):

sin x when A = 1, w=1 and phi = 0 cos x when A=1, w=1 and phi = pi/2

OpenStudy (anonymous):

sin: Φ = 0 cos: Φ = π/2 Gotcha!

hartnn (hartnn):

phi = -pi/2 will also give you a cos function using, \(\sin(\pi/2-x)=\cos x\)